OpenStudy (pottersheep):

Differentiation of y = (e^[x+3])^1/2 Can someone please guide me through?

OpenStudy (phi):

(a^b)^c is the same as a^(bc) so you can write this as \[ y = e^{\frac{x+3}{2}} \]

OpenStudy (phi):

use the rule \[ d e^u = e^u du \]